OpenStudy (anonymous):

The area of a rectangle is 20 meters. How long is the length if the width is 4 meters? 5 16 24 80 A parallelogram is a 4-sided shape formed by one pair of parallel lines. True or False The area of a triangle is 75 square feet. If the height is 5 feet longer than the base, find the dimension of the height. 10 15 20 25

OpenStudy (nincompoop):

AREA of a rectangle = A = L*W 20 = L*W 20 = L*4 solve for L

OpenStudy (anonymous):

=5

OpenStudy (nincompoop):

then that is your answer

OpenStudy (nincompoop):

do you know what a parallelogram looks like?

OpenStudy (anonymous):

no do u mind showing me?

OpenStudy (nincompoop):

use a search engine online

OpenStudy (anonymous):

ok

OpenStudy (nincompoop):

the side of the heights are parallel and the side of the base and top are also parallel 

OpenStudy (anonymous):

so it is true

OpenStudy (nincompoop):

you tell me I basically gave you everything you know

OpenStudy (nincompoop):

let us move on to triangles Area of triangle = \(A_T = \frac{1}{2} (base \times height) \) it says that the height is 5 ft longer than the base this gives us Height = Base + 5 let us try to create our equation \(A_T = 75 = \frac{1}{2} [base \times (base + 5)] \) notice that it seems we are not solving directly for the height, but determining what base is first before we can determine what height i

OpenStudy (nincompoop):

I'd like for you to solve for the base

OpenStudy (anonymous):

ok one second

OpenStudy (anonymous):

the base =15

OpenStudy (nincompoop):

show me your solution please

OpenStudy (nincompoop):

assuming that base is 15 we determined based on the question that, \(\sf Height = Base + 5 \rightarrow Height = 15 + 5 \)

OpenStudy (nincompoop):

aren't these problems easy like doodling?

OpenStudy (anonymous):

yes its just i get confused at some points

OpenStudy (nincompoop):

do not forget your units ft, meters, etc

OpenStudy (anonymous):

so the final answerr of this equation is 25

OpenStudy (nincompoop):

keep working on things that confuse you so you can shrug it off like a cobweb

OpenStudy (nincompoop):

errrr H = 15 + 5 = 25?

OpenStudy (anonymous):

yup

OpenStudy (nincompoop):

I am about to take the medal I gave you back

OpenStudy (anonymous):

lol opps 20

OpenStudy (nincompoop):

FOCUS!

OpenStudy (anonymous):

its like 3 in the morning my brain isnt working xD

OpenStudy (nincompoop):

then go to sleep and work on the problems when your brain IS working

OpenStudy (anonymous):

lol will do

OpenStudy (anonymous):

ty for ur help

OpenStudy (anonymous):

The answer for the last equation was 15 not 20...